Hailstones fall vertically on the bonnet of a car on a horizontal strech of a road the cars bonnet can wisrand impulse of 0.48N.s without sustaining a permanent dent calculate the max mass of the hailstone that will leave the dent in the bonnet of the car it it falls at 8m/s and rebounds at v6m/s after collision lasting 0.002s


1
Expert's answer
2021-03-29T18:54:30-0400
"J=\Delta p=m\Delta v,""m=\dfrac{J}{\Delta v}=\dfrac{J}{v_f-v_i}=\dfrac{0.48\ N\cdot s}{6\ \dfrac{m}{s}-(-8\ \dfrac{m}{s})}=0.034\ kg."